OpenStudy (anonymous):

Using the ratio of perfect squares method, what is square root of 14 rounded to the nearest hundredth?

OpenStudy (anonymous):

@amistre64 could you help me?

OpenStudy (anonymous):

@dtan5457

OpenStudy (amistre64):

ive never come across that method ... srry

OpenStudy (anonymous):

oh. its alright.

geerky42 (geerky42):

Well, \(\sqrt{14} = \sqrt{2\cdot7} = \sqrt2\sqrt7\).

OpenStudy (anonymous):

so i multiply 2 and 7?

geerky42 (geerky42):

multiply \(\sqrt2\) and \(\sqrt7\)

OpenStudy (dumbcow):

i am intrigued... so the goal is to find 2 num a,b such that a^2/b^2 = 14, then sqrt(14) = a/b

geerky42 (geerky42):

Are you familiar with "ratio of perfect squares method?" @dumbcow

OpenStudy (dumbcow):

no im not, im just speculating

OpenStudy (anonymous):

nupe. not one bit. I was previously in pre-al then i trnsferred school now i am suddenly in Algebra

geerky42 (geerky42):

well from quick looks on google, it seems that you just simplify radical, then calculate from there, I guess.

OpenStudy (anonymous):

i got 14 from 2 and 7

OpenStudy (triciaal):

what is the ratio of perfect squares method? is there another name for it ?

geerky42 (geerky42):

I said multiply \(\sqrt2\) and \(\sqrt7\)

OpenStudy (anonymous):

i know. i did. i got \[\sqrt{2}\]

OpenStudy (anonymous):

i mean 14

geerky42 (geerky42):

So \(\sqrt2\times\sqrt7 = 14\)?

OpenStudy (anonymous):

well \[\sqrt{14}\]

geerky42 (geerky42):

well, we are back to where we start. Just convert \(\sqrt2\) and \(\sqrt7\) into decimal, then multiply them together.

OpenStudy (anonymous):

ok

OpenStudy (dumbcow):

@pikapow , just put sqrt(14) in your calculator then round you cant factor any perfect squares out of 14

OpenStudy (anonymous):

3.7224

OpenStudy (anonymous):

right?

geerky42 (geerky42):

If you don't have to show work, there is no point in doing that, so just do what @dumbcow said lol

OpenStudy (anonymous):

was i right though?

geerky42 (geerky42):

Well, I got 3.7416

OpenStudy (anonymous):

me too

OpenStudy (anonymous):

wait. no i didnt

geerky42 (geerky42):

Okay, so round it to nearest hundredth and you have 3.74

geerky42 (geerky42):

Calculate \(\sqrt{14}\)?

OpenStudy (triciaal):